The work function (Φ) for a metal is 7.80×10-19 J. What is the longest wavelength of electromagnetic radiation that can eject an electron from the surface of a piece of the metal?
Answer #1

Longest wavelength = hc / Work function .

Work function =7.80×10-19J

h= plank's constant =6.626×10-34Js

c=velocity of light in vacuum =3×108m/s

Longest wavelength =( 6.626×10-34Js ×3×108m/s)/7.80×10-19

= 2.54×10-7m.​​​​​​​​​ =254nm.